# Exploring Parent-Child Perceptions on Safety in Generative AI: Concerns, Mitigation Strategies, and Design Implications

### v1

@@ 1 Introduction
Generative artificial intelligence (GAI) has swiftly become a vital component of the digital ecosystem, significantly transforming user interactions with technology and opening up new avenues for creativity and innovation. Examples like ChatGPT, which enhances content creation with advanced text generation, and DALL-E, which creates realistic images from text descriptions, illustrate GAI’s impact on digital media and design. This widespread adoption is particularly evident among younger age groups. A study by the UK’s communications regulator Ofcom finds that 79% of online 13-17 year-olds and 40% of 7-12 year-olds in the UK are using generative AI tools and services [1]. As young generations rapidly adopt emerging technologies, it is crucial to consider the potential risks and harms these technologies may bring, including concerns related to discrimination, privacy, and safety. Particularly, the usage of GAI by teens and young adults is closely linked to their mental health [2]. Therefore, it is imperative to thoroughly understand their usage patterns and evaluate the current protective measures. By doing so, we can develop and implement more effective precautions to safeguard their well-being as they engage with these technologies.
Previous studies have examined the risk perceptions and controls of parents across various technologies and applications, including Virtual Reality [3], IoT devices [4], social media platforms [5], and gaming [6]. These studies identified both common and unique risks in various contexts, including explicit content, addiction, cyberbullying, and harassment [7, 8, 9, 10]. Subsequent research has also highlighted parents’ mediation strategies, such as monitoring, restriction, and active mediation [11]. Additionally, studies have pointed out the trade-offs that exist between parental control and children’s autonomy. Since children can benefit from low-risk experiences to develop their ability to identify and manage risks independently in the future, defining the boundaries of parental control and ensuring the right balance remain challenging in children’s risk management [12, 13]. The motivation for our research stems from the need to better understand both parents’ and children’s safety perceptions and coping strategies within the context of Generative AI (GAI), an emerging technology with unique risks and opportunities. We focus on three research questions:
  • RQ1: How did children use GAI and their risk experiences?
@@
Findings. We conducted reddit data analysis and semi-structured interview with seven children and 13 parents to answer these research questions. For RQ1, our findings reveal that child participants use Generative AI (GAI) in diverse and unexpected ways, far beyond the educational or experimental purposes that parent participants assume. Interviews with child participants and analysis of data from Reddit discussions indicate that teenagers often integrate GAI into their daily routines. They use GAI as an friend or confidant, providing emotional support and companionship. Additionally, teenagers reported using GAI to enhance their social interactions, such as incorporating bots into group chats to facilitate conversations and simulate social dynamics. Our research identified a primary reason for this gap: the diverse sources from which children are informed and introduced to AI. Besides their parents, children learn about GAI through influencer videos, targeted advertisements on social media, and peer recommendations. These sources often introduce them to new and experimental GAI platforms, which typically carry more risks compared to more established ones. Furthermore, both parents and children hold significant misconceptions about GAI’s data collection, sharing, and processing practices.
@@
For RQ2, we uncovered significant security and privacy concerns associated with children’s use of generative artificial intelligence (GAI), which parents frequently overlook due to unawareness of the varied interactions children have with these technologies. Teenagers expressed specific worries about GAI in social contexts. They were concerned about becoming overly dependent on virtual relationships with chatbots, which could exacerbate their challenges in forming real-life social connections. Additionally, they noted the potential misuse of GAI in group chats to create distressing or harassing content. Moreover, child participants shared concerns regarding the unauthorized use of their personal data to create GAI chatbots or memes, and the risk of strangers misusing their online images to glean personal information. They also encountered NSFW chatbots promoted by peers and social media, and some even created chatbots to spread harmful content, such as racist remarks. Parents, however, often underestimated these risks. While they were aware of GAI collecting basic information like names and locations, they did not fully appreciate the extent of sensitive data their children might share with GAI perceived as friends or confidants. This data could include details of personal traumas, medical records, and private aspects of their social and sexual lives.
@@
Contributions. This paper offers three main contributions to the field. First, we identified unique usage patterns and associated risks of GAI among teenagers. Second, we investigated the existing GAI mediation strategies employed by parents and identified challenges they face. Third, we evaluated both children’s and parents’ perspectives on desired GAI mediation and derived constructive design implications for future GAI systems.
@@ 3.1.1 Data Collection
To comprehensively cover content related to our research questions on teenagers and Generative AI, we first created a list of search keywords by identifying close terminologies related to “teenager” and “Generative AI”. We then followed specific criteria to select subreddits: these subreddits should focus either on the teenager community or Generative AI technology. We chose subreddits with the most active users online during our browsing sessions. The full list of subreddits and search keywords used is detailed in Table III.
We first performed open searches combining these keywords across the entire Reddit platform. Then, we searched for all terms related to “teenager” within Generative AI subreddits (r/midjourney, r/ChatGPT, r/OpenAI), and all terms related to “Generative AI” within teenager subreddits (r/teenagers, r/AskTeenGirls, r/askteenboys, r/BisexualTeens). Finally, we removed any duplicates from the search results. The number of posts collected from each subreddit is detailed in Table III.
@@ 3.2 Interview Study
In addition to our Reddit study, we conducted a semi-structured interview study between January and May 2024. We interviewed 7 children and 13 parents to gain a deeper understanding of their experiences and risk perceptions related to Generalized Artificial Intelligence (GAI). We also aimed to identify unmet needs for parental control features in GAI tools. We developed two tailored versions of a user interview protocol, one for parents and one for children, aligned with our research questions. The protocol was structured according to the following topical sections: (1) current practice-related parental mediation strategies to ensure privacy, security, and safety for teenagers; (2) current use of GAI tools by both parents and children and their attitudes towards the risk of GAI tools (3) existing parental control framework or design fits the GAI tool scenario in protecting children’s security, privacy and safety. The questions in the children’s interview protocol were tailored to accommodate different age groups, ensuring developmental appropriateness and clarity, and thereby enabling us to gather rich and meaningful insights from both children and parents. These interviews were typically around 1 hour in length. We have included the interview protocol in Appendix 8.1.
@@ 3.2.1 Participant Recruitment
Our eligibility criteria and method were broad in order to capture a wide spectrum of experience with GAI application, gender, and age (13-17). Participants for the interviews were recruited through several methods: (i) To recruit teenagers, we sent out a student survey to a local public high school in the US. If students expressed interest in participating, we obtained written informed consent from parents via email, following our institution’s approved protocols and procedures for securing consent; (ii) We also recruited parents through Prolific by distributing a parent survey, using the following screening criteria: fluency in English and having at least one child aged 13-17 years old. Parents who expressed interest in participating will be interviewed first; then, we will provide the opportunity to invite their children to participate in our study as wells. If parents expressed interest in having their children participate in the study and provided informed consent, we shared the student survey with them via email. All participants who completed the interviews were compensated with a \$25 Amazon gift card. The demographics of our participants are summarized in Table II in the Appendix.
@@ 3.2.2 Interview Procedure
We emailed parent participants one day before their scheduled interview to obtain their written consent. Similarly, for teenagers, we sent emails to both the teenagers and their parents to secure consent from both parties. We conducted semi-structured interviews with each participant via Zoom, where we obtained verbal consent from all participants before recording the sessions. Each interview was audio recorded and transcribed verbatim. To address our research questions, each interview was divided into three main sections: initially, we discussed family background and the parental controls in place; secondly, we explored participants’ experiences with GAI and their perceptions of associated risks; finally, we delved into their strategies and challenges in mediating GAI usage. Due to space constraints, the detailed interview procedure and the complete list of interview questions are included in the Appendix 8.1.
@@ 4.1 Overview
In total, we analyzed 181 posts/comments. In each subreddit, we filtered through comments/posts related to both teenagers and generative AI. Most of the data (98%) collected from our Reddit search were not relevant to our research questions, as they were often not posted by teenagers or did not pertain to the topic of generative AI. The detailed list of analyst subreddits (2%), along with the number of collected and analyzed posts/comments, is provided in Table III. Through the thematic coding, we found most posts/comments are under two categories: Teenager GAI Usage Patterns and Teenager GAI Concerns. We provided the frequency of each high-level theme in Table I. Our findings revealed that teenagers are using GAI widely and diversely in their daily lives, encompassing emotional support, social interactions, education, entertainment, and various risk-seeking behaviors. They also express a broad spectrum of concerns about GAI online, ranging from societal harms such as replacing human labor and intellectual property infringement to security and privacy concerns, including the misuse of private information with GAI technology. Our user interview findings overlap with and complement these two topics, so we have combined them. We use the term “users” to represent findings from Reddit and “participants” to report findings from user interviews. Detailed findings on teenagers’ GAI usage and concerns are presented in the remaining sections.

### v2

@@ 1 Introduction
Generative artificial intelligence (GAI) has swiftly become a vital component of the digital ecosystem, significantly transforming user interactions with technology and opening up new avenues for creativity and innovation. Examples like ChatGPT, which enhances content creation with advanced text generation, and DALL-E, which creates realistic images from text descriptions, illustrate GAI’s impact on digital media and design. In contrast, Character.ai offers a unique experience by enabling users to engage with character-based AI chatbots, allowing for more personalized and interactive conversations, often mimicking human-like interactions. This widespread adoption is particularly evident among younger age groups. A study by the UK’s communications regulator Ofcom finds that 79% of online 13-17 year-olds and 40% of 7-12 year-olds in the UK are using generative AI tools and services [1]. As young generations rapidly adopt emerging technologies, it is crucial to consider the potential risks and harms these technologies may bring, including concerns related to privacy, and safety. Particularly, the usage of GAI by teens and young adults is closely linked to their mental health [2]. Therefore, it is imperative to thoroughly understand their usage patterns and evaluate the current protective measures. By doing so, we can develop and implement more effective precautions to safeguard their well-being as they engage with these technologies.
Previous studies have examined the risk perceptions and parental controls across various technologies and applications, including Virtual Reality [3], IoT devices [4], social media platforms [5], and gaming [6]. These studies identified both common and unique risks in various contexts, including explicit content, addiction, cyberbullying, and harassment [7, 8, 9, 10]. Subsequent research has also highlighted parents’ mediation strategies, such as monitoring, restriction, and active mediation [11]. Additionally, studies have pointed out the trade-offs that exist between parental control and children’s autonomy. Since children can benefit from low-risk experiences to develop their ability to identify and manage risks independently in the future, defining the boundaries of parental control and ensuring the right balance remain challenging in children’s risk management [12, 13]. The motivation for our research stems from the need to better understand both parents’ and children’s safety perceptions and coping strategies within the context of Generative AI (GAI), an emerging technology with unique risks and opportunities. We focus on three research questions:
  • RQ1: How did teenagers use GAI?
@@
To address these questions, we utilized a mixed-methods approach that combined Reddit content analysis, followed by semi-structured interviews involving teenagers and parents. Initially, we collected 712 posts and 8,533 comments from relevant subreddits, providing naturalistic insights into how teenagers engage with GAI. This was followed by semi-structured interviews with seven teenagers and 13 parents, designed to explore their in-depth perspectives on GAI use, safety concerns, privacy issues, and the strategies parents employ to mediate their teenagers’ interactions with these technologies.
@@
Findings. For RQ1, our findings reveal that child participants use Generative AI (GAI) in diverse and unexpected ways, far beyond the educational or experimental purposes that parent participants assume. Interviews with child participants and analysis of data from Reddit discussions indicate that teenagers often integrate GAI into their daily routines. They use GAI as an friend or confidant, providing emotional support and companionship. Additionally, teenagers reported using GAI to enhance their social interactions, such as incorporating bots into group chats to facilitate conversations and simulate social dynamics. Our research identified a primary reason for this gap between parents and children: the diverse sources from which children are informed and introduced to AI. Besides their parents, children learn about GAI through influencer videos, targeted advertisements on social media, and peer recommendations. These sources often introduce them to new and experimental GAI platforms, which typically carry more risks compared to more established ones. Furthermore, both parents and children hold significant misconceptions about GAI’s data collection, sharing, and processing practices. Teenagers often view GAI as either functioning like a search engine or a vast database, believing it directly pulls responses from the internet or provides answers from a stored set of responses. Similarly, many parents shared this search engine misconception, but some parents also incorrectly assumed that GAI platforms perform fact-checking and verification before generating responses.
@@
For RQ2, we uncovered significant security and privacy concerns associated with children’s use of generative artificial intelligence (GAI), which parents frequently overlook due to unawareness of the varied interactions children have with these technologies. Child participants expressed specific worries about GAI in social contexts. They were concerned about becoming overly dependent on virtual relationships with chatbots, which could exacerbate their challenges in forming real-life social connections. Additionally, they noted the potential misuse of GAI in group chats to create distressing or harassing content. Moreover, child participants shared concerns regarding the unauthorized use of their personal data to create GAI chatbots or memes, and the risk of strangers misusing their online images to glean personal information. They also encountered not safe for work (NSFW) chatbots promoted by peers and social media, and some even created chatbots to spread harmful content, such as racist remarks. Parents, however, often underestimated these risks. While they were aware of GAI collecting basic information like names and locations, they did not fully appreciate the extent of sensitive data their children might share with GAI perceived as friends or confidants. This data could include details of personal traumas, medical records, and private aspects of their social and sexual lives.
@@
Contributions. This paper offers three main contributions to the field. First, we identified unique usage patterns and associated risks of GAI among teenagers. Second, we investigated the existing GAI mediation strategies employed by parents and identified challenges they face. Third, we evaluated both children’s and parents’ perspectives on desired GAI mediation and derived constructive design implications for future GAI systems. Specifically, we recommend that GAI platforms incorporate age-appropriate content moderation, customizable parental control tools, and transparent risk disclosures to better address the safety needs of teenage users.
@@ 3.1.1 Data Collection
To comprehensively cover content related to our research questions on teenagers and Generative AI, we first created a list of search keywords by identifying close terminologies related to “teenager” (general keywords) and “Generative AI” (technology-focused keywords). We utilized a combination of general and technology-focused keywords in our search. We employed general terms such as teen, teenager, adolescent, and high school student. These keywords were designed to capture posts authored by or discussing teenagers. For the technology focus, we used terms such as Generative AI, AI chatbot, artificial intelligence, AI, ChatGPT, DALL-E, and Midjourney. These keywords targeted discussions specifically about the use of popular AI tools and platforms. We conducted open searches combining these keywords across Reddit to gather data from various subreddits. Other than open searches, we also applied specific criteria to select subreddits, ensuring comprehensive coverage of relevant discussions: these subreddits should focus either on the teenager community or Generative AI technology. We chose subreddits with the most active users online during our browsing sessions. The full list of subreddits and search keywords used is detailed in Table I.
We first performed open searches combining these keywords across the entire Reddit platform. Then, we searched for all terms related to “teenager” within Generative AI subreddits (r/midjourney, r/ChatGPT, r/OpenAI), and all terms related to “Generative AI” within teenager subreddits (r/teenagers, r/AskTeenGirls, r/askteenboys, r/BisexualTeens). Finally, we removed any duplicates from the search results. The number of posts collected from each subreddit is detailed in Table I.
@@ 3.1.2 Analysis
Subreddit # of Posts Pulled # of Comments Pulled # of Related Posts/Comments
Open search
all subreddits
159 2034 32
teenager r/teenagers 243 4294 90
r/BisexualTeens 45 203 24
r/askteenboys 10 18 7
r/AskTeenGirls 7 15 11
GAI r/ChatGPT 158 1783 10
r/OpenAI 9 34 7
r/midjourney 81 152 0
TABLE I: Overview of statistics in the Reddit dataset.
@@ 3.2 Interview Study
In addition to our Reddit study, we conducted a semi-structured interview study between January and May 2024. We interviewed 7 children and 13 parents to gain a deeper understanding of their experiences and risk perceptions related to Generalized Artificial Intelligence (GAI). We also aimed to identify unmet needs for parental control features in GAI tools. We developed two tailored versions of a user interview protocol, one for parents and one for children, aligned with our research questions. The protocol was structured according to the following topical sections: (1) current practice-related parental mediation strategies to ensure privacy, security, and safety for teenagers; (2) current use of GAI tools by both parents and children and their perceptions towards the risk of GAI tools (3) existing parental control framework or design fits the GAI tool scenario in protecting children’s security, privacy and safety. The questions in the children’s interview protocol were tailored to accommodate different age groups, ensuring developmental appropriateness and clarity, and thereby enabling us to gather rich and meaningful insights from both children and parents. These interviews were typically around one hour in length.
@@ 3.2.1 Participant Recruitment
We recruited parent participants through Prolific based on the following inclusion criteria: (1) prior experience with Generative AI, including the duration of use and the specific platforms they have engaged with , (2) residing in the U.S. and fluent in English, and (3) having at least one child aged 13-17. Teenager participants were recruited from local high schools and through Prolific, with some parents introducing their own children as well. This approach ensured diversity in our participant pool, with variations in both the platforms used and the duration of their Generative AI experience.
(i) To recruit teenagers, we sent out a student survey to a local public high school in the US. If students expressed interest in participating, we obtained written informed consent from parents via email, following our institution’s approved protocols and procedures for securing consent; (ii) We also recruited teenagers through Prolific by distributing a parent survey. Parents who expressed interest in participating will be interviewed first; then, we will provide the opportunity to invite their children to participate in our study as wells. If parents expressed interest in having their children participate in the study and provided informed consent, we shared the student survey with them via email. All participants who completed the interviews were compensated with a \$25 Amazon gift card. The demographics of our participants are summarized in Table II.
ID Age Gender Used GAI Months of Use State
P1 17 (teen) Male ChatGPT 1 to 3 months IL
P2 14 (teen) Male ChatGPT, Bard, Character.ai 1 to 3 months IL
P3 16 (teen) Male ChatGPT, CoPilot, Logo ai generator 6 months to 1 year IL
P4 13 (teen) Female ChatGPT 1 to 3 months NY
P5 15 (teen) Male ChatGPT, Midjourney, Deepfake Voice 6 months to 1 year IL
P6 13 (teen) Male ChatGPT, Character.ai Less than 1 month OH
P7 15 (teen) Male DALL-E, Character.ai, Chai AI 6 months to 1 year VA
P8 54 Male ChatGPT 3 to 6 months DC
P9 37 Male ChatGPT 3 to 6 months CO
P10 34 Female ChatGPT 6 months to 1 year TX
P11 52 Female ChatGPT 1 to 3 months NY
P12 44 Male ChatGPT 3 to 6 months OH
P13 35 Female ChatGPT 3 to 6 months MI
P14 49 Female ChatGPT, DALL-E, Midjourney 6 months to 1 year VA
P15 35 Female ChatGPT 1 to 3 months VA
P16 37 Female ChatGPT 3 to 6 months SC
P17 36 Male ChatGPT, DALL-E, Midjourney More than 1 year IN
P18 42 Male ChatGPT, DALL-E, Midjourney 6 months to 1 year NV
P19 43 Female ChatGPT, DALL-E, Diffit, Magic School AI 3 to 6 months GA
P20 35 Female ChatGPT, DALL-E 3 to 6 months FL
TABLE II: Participants demographics with usage patterns and locations highlighted.
@@ 3.2.2 Interview Procedure
We emailed parent participants one day before their scheduled interview to obtain their written consent. Similarly, for teenagers, we sent emails to both the teenagers and their parents to secure consent from both parties. We conducted semi-structured interviews with each participant via Zoom, where we obtained verbal consent from all participants before recording the sessions. Each interview was audio recorded and transcribed verbatim. To address our research questions, each interview was divided into three main sections: initially, we discussed family background and the parental controls in place; secondly, we explored participants’ experiences with GAI and their perceptions of associated risks; finally, we delved into their strategies and challenges in mediating GAI usage.
3.2.3 Interview protocol
Interview with Teenagers. For the interviews with teenagers, the first section focused on gathering background information about their living environment, including the number of people in their household, device access, and the type of devices they use. This helped us understand if they share devices and with whom. We asked participants, “Do you personally own these devices or do you share them with other family members? If you share, can you specify with whom?” We then inquired about any parental controls implemented on their devices or technologies, asking, “Have they set up any rules or restrictions for your device/technology usage, such as for gaming, social media, or any apps/websites you use?” We also asked about how often their parents monitor the devices and how well they understand these functionalities. Finally, we asked about their feelings regarding the current parental controls.
In the next section, we focused on their current usage and understanding of generative AI (GAI) tools. We started by discussing their specific use of GAI and what they like about these applications. We encouraged them to imagine and describe how GAI works, for example, by asking, “Can you create a drawing or describe how you imagine tools like ChatGPT work?” This helped us gauge their level of understanding. We also explored how they would explain GAI to others, to understand their perception and how they communicate about this technology. We asked if their parents taught them how to use GAI and the extent of discussions they had regarding GAI’s safety, benefits, and risks, with questions like, “Have your parents had discussions with you about the safety and risks associated with using GAI tools?” We concluded this section by asking if they believe there are potential risks parents should be aware of for their children.
In the final section, we asked about any rules provided by both their parents and teachers when using GAI. We inquired if they think having rules impacts their usage of GAI. We then asked for their expectations of control for GAI and if they have any further concerns related to generative AI, such as, What kind of rules or controls do you think would be helpful for kids using GAI tools?”
Interview with Parents. Like the interviews with teenagers, the parents’ interview protocol aimed to address our research questions. We began by discussing their current practices of parental controls and their reasons for using them. We then asked about the effectiveness of these controls and gauged their perception of how their children feel about these controls, including any conflicts that might have arisen due to their use.
Next, we explored their familiarity with generative AI (GAI) tools, their understanding of how these tools work, and how they would explain GAI tools to others. We encouraged them to draw their understanding of how GAI tools function, including data flow, and asked, “What data do you think generative AI tools collect from users like you?” We also inquired about their feelings regarding the data processing of GAI tools and their children’s usage of these tools. We asked about their current parental controls for GAI tools, if any, with follow up questions like, “Can you describe how you monitor your child’s use of these devices?” We then discussed their concerns about their children using GAI tools and whether they had any discussions with their children about these concerns.
Finally, we asked parents how they would respond to their child using new GAI tools, including setting up rules or restrictions for their children’s GAI tool usage and whether these would be helpful. We inquired if they were aware of any existing tools for GAI parental control and how they would like to set up rules or restrictions. We then asked how their children might feel about these expected parental controls. We showed them some features for parental control of GAI and asked for their opinions to inform future design. Example features included content filtering and session duration limits. To understand their priorities, we concluded with, “Of the features mentioned, which ones do you think would be the most critical for your child’s safety and why?”
The full list of interview questions is available at https://github.com/SPresearch/Interview-protocol.
@@ 3.2.4 Interview Data Analysis
3.2.5 Limitation
While our study incorporated well-known GAI platform subreddits such as ChatGPT, there are inherent limitations in our data collection process due to the rapidly evolving nature of GAI technology. For instance, although Character.ai emerged as highly relevant during participant interviews and Reddit discussions, it was not prominently featured during the initial stages of our Reddit data collection. This indicates that other similarly popular or emerging platforms might have also been overlooked. Expanding the scope of included platforms in future research will ensure a more comprehensive representation of GAI usage among teenagers, capturing the full spectrum of both well-established and emerging tools.
A potential limitation of our study is the use of ChatGPT as an example in the interview guide, which may have primed participants to focus on this platform. ChatGPT was mentioned due to its widespread recognition, ensuring participants understood the concept of Generative AI. Future studies could consider either not mentioning any specific GAI tools or mentioning a wide range of GAI tools as examples.
3.2.6 Ethics & Data Protection
Before participating in the interview study, participants provided their consent for both the study and audio recording through a consent form by email. In addition, we require parental consent in both verbal format via Zoom and written format through email before each child participant interview. Participants were informed of their right to withdraw from the study at any time without repercussions or loss of benefits, and they had the option to skip any questions during the interview. We assured them that their quotes would be used in a non-identifiable manner. This approach, while limiting the quotes or descriptions we can report, allowed participants to speak freely about their experiences. Additionally, we addressed any questions participants had about the procedure and purpose of the study and provided a debriefing after the interviews.
Transcripts were pseudonymized and stored in a secure university cloud environment for storage and collaborative coding. Each participant received \$25 after interview, based on an expected interview duration of up to 60 minutes, equivalent to a rate of \$21.24 per hour [47], the average rate in USA. The study was approved by our institution’s ethical review board and data protection office.
@@ 4.1 Overview
In total, we analyzed 181 posts/comments. In each subreddit, we filtered through comments/posts related to both teenagers and generative AI. Most of the data (98%) collected from our Reddit search were not relevant to our research questions, as they were often not posted by teenagers or did not pertain to the topic of generative AI. The detailed list of analyst subreddits (2%), along with the number of collected and analyzed posts/comments, is provided in Table I. Through the thematic coding, we found most posts/comments are under two categories: Teenager GAI Usage Patterns and Teenager GAI Concerns. Our findings revealed that teenagers are using GAI widely and diversely in their daily lives, encompassing emotional support, social interactions, education, entertainment, and various risk-seeking behaviors. They also express a broad spectrum of concerns about GAI online, ranging from societal harms such as replacing human labor and intellectual property infringement to security and privacy concerns, including the misuse of private information with GAI technology. Our user interview findings overlap with and complement these two topics, so we have combined them. We use the term “users” to represent findings from Reddit and “participants” to report findings from user interviews. Detailed findings on teenagers’ GAI usage and concerns are presented in the remaining sections. All published quotes are paraphrased from existing non-deleted posts to preserve pseudonymity.